# Mitigating Lost-in-Retrieval Problems in Retrieval Augmented Multi-Hop Question Answering

### v1

@@ 1 Introduction
Large language models (LLMs) OpenAI (2023); Zeng et al. (2024); Yang et al. (2024); Li et al. (2024a) have exhibited promising performance on a wide range of natural language processing tasks, such as machine translation Zhu et al. (2023), text summarization Wu et al. (2021), question answering (QA) Daull et al. (2023). While LLMs possess strong reasoning abilities, they still face challenges such as outdated knowledge and lack of domain-specific expertise Mousavi et al. (2024), which lead to incorrect outputs Xu et al. (2024b).
To fill the gap between LLMs’ memory and real-world knowledge, retrieval-augmented generation (RAG) Lewis et al. (2020); Gao et al. (2023) is widely used to retrieve the knowledge that is relevant to the user’s question to improve the LLMs’ QA performance. When answering multi-hop questions, LLMs generally utilize a question decomposition strategy in the chain-of-thought outputs, which decomposes the input question into multiple simpler sub-questions. However, in this decomposition process, we find that when a sub-question lacks a clear entity and instead uses demonstrative pronouns, the retrieval performance drops sharply. We refer to this phenomenon as “lost-in-retrieval”.
Figure 1 presents a real-world example in solving a multi-hop question using RAG combined with question decomposition. In this example, the second sub-question, “What was the home city of this author?”, lacks a clear entity of the author. As a result, it leads to retrieval errors, which ultimately cause the final answer to be incorrect. The right answer is Vienna. To give the correct answer, we need to identify the specific key entities in the sub-questions so as to improve the retrieval performance.
@@ 2.2 Multi-hop QA
Multi-hop QA is an ideal scenario for evaluating RAG systems, since it requires strong capabilities from both the knowledge retriever and the answer generator. RoG Luo et al. (2024) adopts a planning-retrieval-reasoning paradigm, using relation paths in KGs to guide the retrieval of effective reasoning paths. EfficientRAG Zhuang et al. (2024) fine-tunes the DeBERTa-v3-large model He et al. (2021) to construct a labeler and a filter for handling multi-round queries, reducing the frequency of LLM calls. OneGen Zhang et al. (2024) unifies generation and retrieval by fine-tuning the model to perform both tasks simultaneously in a single-step inference. Many existing multi-hop QA methods use LLMs for query decomposition Gao et al. (2023). However, as demonstrated in the empirical study in Section 1, this strategy suffers from “lost-in-retrieval”. Our work resolves this issue without fine-tuning and frequent API calls.
@@ 3.1 Sentence Graph with Entity Indexing
In our method, the completion of missing key entities in sub-questions is a crucial step. To facilitate entity completion, it is essential to identify all named entities retrievable from the given texts. Therefore, for efficiency consideration, we first extract named entities from texts $s_{i}$ using spaCy, resulting in an entity set $\mathcal{E}_{i}$. In this process, we also store the mappings between each entity and all its sentences. Then, to conveniently obtain all the information of an entity from scattered texts for the following knowledge retrieval, we propose to construct a sentence graph with named entities as edge labels, where each node represents a sentence and the edge between nodes indicates that the two sentences describe the same entity. We denote the node (i.e., sentence) set as $C=\{s_{1},s_{2},\dots,s_{n}\}$, which has already been obtained in the previous entity extract process. As for the edge, the entity-sentence mappings can be used to mine edges. However, relying solely on entity co-occurrence edges is insufficient for effective knowledge retrieval. Besides, the sentence-level retrieval is too fine-grained. We should enhance the associations between sentences for border and comprehensive retrieval. Finally, we consider the following three types of edges in our sentence graph:
@@ 3.2 Sentence and Entity Retrieval
Given a sub-question, we first calculate its embedding similarity (e.g., inner product) with all sentences in the sentence graph. This can be done quickly by matrix multiplication. We then filter our the sentences with low similarity to narrow down the retrieval candidates. Next, we use a cross-encoder to assess the relevance of each candidate sentence within the context of the sub-question, and finally, we select the top-$k$ sentences as seed sentences.
@@ 3.3 Sub-question Rewriting
As we have mentioned and validated in Sect. 1 that the retrieval performance degrades if sub-questions lack necessary named entities. To resolve this problem, we propose to rewrite sub-questions. First, we determine if a sub-question needs rewriting by checking for the presence of pronouns (such as “this”, “it”, “they”, etc.). If these pronouns exist, we feed both the current sub-question and the previous sub-questions along with their answers into an LLM to rewrite the current sub-question. When the previous sub-questions were not adequately answered, we cannot rewrite the sub-questions. In this case, we summarize their corresponding context and incorporate the summary into the context of the current sub-question. Our sub-question rewriting method serves two purposes. First, it mitigates the degradation in retrieval performance caused by missing entities. Second, since the previous sub-question is closely related to the current one, preserving its key information supports the reasoning process of the current sub-question.
@@ 3.4 Answer and Context Integration
This method generates the answer to the original question by utilizing each sub-question and its answer. It relies solely on the information from the sub-questions’ answers, without external interference. Since the decomposition of sub-questions can be regarded as a reasoning process, this method enables the LLM to infer the original question’s answer. Moreover, this method ensures that the LLM processes the relevant text of only one sub-question at a time, avoiding performance degradation caused by the LLM’s weak long-context processing capabilities when processing multiple sub-questions simultaneously. However, if a sub-question is answered incorrectly or left unanswered, it can significantly impact the final result. To reduce such impact, we consider the following method that integrates all retrieved sentences to enrich the context.
@@
We remove duplicate contexts from all retrieved sentences (without using sub-questions) and use a cross-encoder to rerank the sentences for generating the final answer. This method is similar to traditional RAG, which uses only relevant text to generate the answer. It helps mitigate the impact of errors in sub-question decomposition or answers by focusing on the retrieved sentences, rather than relying solely on the sub-question answers. However, this method requires the LLM to have strong long-context processing capabilities. Compared to answering each sub-question individually, the merged context may contain more noisy information, which can negatively impact the final answer.
@@ 4 Experiments
In this section, we report the experimental results and analysis to evaluate the effectiveness and efficiency of our method for multi-hop QA.
@@ 4.4 Retrieval Performance Analysis
To validate whether our method effectively addresses the “lost-in-retrieval” problems, we conduct additional retrieval experiments. Since our method divides the context into sentences, which differs rom NaiveRAG in granularity, we standardize the chunk size in ChainRAG to match NaiveRAG for fair comparison. From Table 2, we observe that after rewriting, the Recall@2 of sub-question 2 shows a marked improvement. Notably, on the MuSiQue and HotpotQA datasets, the Recall@2 of the rewritten sub-question 2 even exceeds that of sub-question 1, as the rewriting process uses information from sub-question 1 to retrieve more relevant content.
@@ 4.7 Case Study
We present an example question from MuSiQue and compare the RAG process of NaiveRAG w/ QD and ChainRAG in Table 3. As NaiveRAG w/ QD does not rewrite the sub-question, its second sub-question lacks a clear entity, leading to retrieval errors and causing the "lost-in-retrieval" problems. This ultimately results in an incorrect answer. In the example, the context comes from the first chunk retrieved, which is also the source of the final answer. For ChainRAG, by using the answer from the first sub-question to complete the entity in the second sub-question, the retrieved information becomes more accurate, on which the LLM then provides the correct answer.
@@ Acknowledgments
This work was funded by National Natural Science Foundation of China (Nos. 62406136 and 62272219), Natural Science Foundation of Jiangsu Province (No. BK20241246), and Postdoctoral Fellowship Program of CPSF (No. GZC20240685).
@@ Appendix C Detailed Results of Ablation Study
Tables 5 and 6 present the complete results of our ablation study. After removing certain processes from ChainRAG, we observe varying degrees of performance degradation in QA tasks. When the rewriting phase is removed, the performance on MuSiQue and HotpotQA drops significantly, while the performance on 2Wiki dataset sees a minor decrease. This is mainly because many of the questions in the 2Wiki dataset, after being decomposed, result in two parallel sub-questions with no dependency, which avoids the “lost-in-retrieval” problems. For MuSiQue, removing the SS edges has the greatest impact on performance, while for 2Wiki, it is the removal of the EC edges that has the most significant effect. In HotpotQA, the SA edges have the largest impact when removed. When we remove the entire graph and implement the indexing process according to NaiveRAG, we observe a noticeable decline in performance, with the CxtInt method being more significantly affected. This indicates the effectiveness of our method, which involves building an index using sentences and entities and performing retrieval on the sentence graph.

### v2

@@ 1 Introduction
Large language models (LLMs) OpenAI (2023); Zeng et al. (2024); Yang et al. (2024); Li et al. (2024a) have exhibited promising performance on a wide range of natural language processing tasks, such as machine translation Zhu et al. (2023), text summarization Wu et al. (2021), question answering (QA) Daull et al. (2023). While LLMs possess strong reasoning abilities, they still face challenges such as outdated knowledge Mousavi et al. (2024) and lack of domain-specific expertise Chen (2024), which lead to incorrect outputs Xu et al. (2024b).
To fill the gap between LLMs’ memory and real-world knowledge, retrieval-augmented generation (RAG) Lewis et al. (2020); Gao et al. (2023); Luo et al. (2023) is widely used to retrieve the knowledge that is relevant to the user’s question to improve the LLMs’ QA performance. Unlike knowledge-enhanced finetuning Luo et al. (2024b), RAG operates without updating LLM parameters. When answering multi-hop questions, RAG-based methods generally utilize a question decomposition strategy to decompose the input question into multiple simpler sub-questions. However, in this decomposition process, we find that when a sub-question lacks a clear entity and instead uses demonstrative pronouns, the retrieval performance drops sharply. We refer to this phenomenon as “lost-in-retrieval”.
Figure 1 presents a real-world example in solving a multi-hop question using RAG combined with question decomposition. In this example, the second sub-question, “What was the home city of this author?”, lacks a clear entity of the author. As a result, it leads to retrieval errors, which ultimately cause the final answer to be incorrect. The right answer is Vienna. To give the correct answer, we need to identify the specific key entities in the sub-questions to improve the retrieval performance.
@@ 2.2 Multi-hop QA
Multi-hop QA is an ideal scenario for evaluating RAG systems, since it requires strong capabilities from both the knowledge retriever and the answer generator. RoG Luo et al. (2024a) adopts a planning-retrieval-reasoning paradigm, using relation paths in KGs to guide the retrieval of effective reasoning paths. EfficientRAG Zhuang et al. (2024) fine-tunes the DeBERTa-v3-large model He et al. (2021) to construct a labeler and a filter for handling multi-round queries, reducing the frequency of LLM calls. OneGen Zhang et al. (2024) unifies generation and retrieval by fine-tuning the model to perform both tasks simultaneously in a single-step inference. Many existing multi-hop QA methods use LLMs for query decomposition Gao et al. (2023). However, as demonstrated in the empirical study in Section 1, this strategy suffers from “lost-in-retrieval”. Our work resolves this issue without fine-tuning and frequent API calls.
@@ 3.1 Sentence Graph with Entity Indexing
In our method, the completion of missing key entities in sub-questions is a crucial step. To facilitate entity completion, it is essential to identify all named entities retrievable from the given texts. Therefore, for efficiency consideration, we first extract named entities from texts $s_{i}$ using spaCy, resulting in an entity set $\mathcal{E}_{i}$. In this process, we also store the mappings between each entity and all its sentences. Then, to conveniently obtain all the information of an entity from scattered texts for the following knowledge retrieval, we propose to construct a sentence graph with named entities as edge labels, where each node represents a sentence and each edge between nodes indicates that the two sentences describe the same entity. We denote the node set as $C=\{s_{1},s_{2},\dots,s_{n}\}$, which has already been obtained in the previous entity extract process. As for the edge, the entity-sentence mappings can be used to mine edges. However, relying solely on entity co-occurrence edges is insufficient for effective knowledge retrieval. Besides, the sentence-level retrieval is too fine-grained. We should enhance the associations between sentences for border and comprehensive retrieval. Finally, we consider the following three types of edges in our sentence graph:
@@ 3.2 Sentence and Entity Retrieval
Given a sub-question, we first calculate its embedding similarity (e.g., inner product) with all sentences in the sentence graph. This can be done quickly by matrix multiplication. We then filter out the sentences with low similarity to narrow down the retrieval candidates. Next, we use a cross-encoder to assess the relevance of each candidate sentence within the context of the sub-question, and finally, we select the top-$k$ sentences as seed sentences.
@@ 3.3 Sub-question Rewriting
As we have mentioned and validated in Sect. 1 that the retrieval performance degrades if sub-questions lack necessary named entities. To resolve this problem, we propose to rewrite sub-questions. First, we determine if a sub-question needs rewriting by checking for the presence of pronouns (such as “this”, “it”, “they”, etc.). If these pronouns exist, we feed both the current sub-question and the previous sub-questions along with their answers into an LLM to rewrite the current sub-question. When the previous sub-questions were not answered, we cannot rewrite the sub-questions. In this case, we summarize their corresponding context and incorporate the summary into the context of the current sub-question. Our sub-question rewriting method serves two purposes. First, it mitigates the degradation in retrieval performance caused by missing entities. Second, since the previous sub-question is closely related to the current one, preserving its key information supports the reasoning process of the current sub-question.
@@ 3.4 Answer and Context Integration
This method generates the answer to the original question by utilizing each sub-question and its answer. It relies solely on the information from the sub-questions’ answers, without external interference. Since the decomposition of sub-questions can be regarded as a reasoning process, this method enables the LLM to infer the original question’s answer. Moreover, this method ensures that the LLM processes the relevant text of only one sub-question at a time, avoiding performance degradation caused by the LLM’s weak long-context processing capabilities when processing multiple sub-questions simultaneously. Consequently, this method does not require the LLM strong long-context processing capabilities, as the final synthesis relies exclusively on the concise information from the sub-questions and their generated answers, without utilizing any retrieved relevant text. However, if a sub-question is answered incorrectly or left unanswered, it can significantly impact the final result. To reduce such impact, we consider the following method that integrates all retrieved sentences to enrich the context.
@@
We remove duplicate contexts from all retrieved sentences (without using sub-questions) and use a cross-encoder to rerank the sentences for generating the final answer. This method is similar to traditional RAG, which uses only relevant text to generate the answer. It helps mitigate the impact of errors in sub-question decomposition or answers by focusing on the retrieved sentences, rather than relying solely on the sub-question answers. However, this method explicitly requires the LLM to have strong long-context processing capabilities to effectively process the merged and potentially extensive context. Compared to answering each sub-question individually, the merged context may contain more noisy information and negatively impact the final answer.
@@ 4 Experiments
In this section, we report the experimental results and analysis to evaluate the effectiveness and efficiency of our method for multi-hop QA. Our source code is available at GitHub. [^1] https://github.com/nju-websoft/ChainRAG
@@ 4.4 Retrieval Performance Analysis
To validate whether our method effectively addresses the “lost-in-retrieval” problems, we conduct additional retrieval experiments. Since our method divides the context into sentences, which differs from NaiveRAG in granularity, we standardize the chunk size in ChainRAG to match NaiveRAG for fair comparison. From Table 2, we observe that after rewriting, the Recall@2 of sub-question 2 shows a marked improvement. Notably, on the MuSiQue and HotpotQA datasets, the Recall@2 of the rewritten sub-question 2 even exceeds that of sub-question 1, as the rewriting process uses information from sub-question 1 to retrieve more relevant content.
@@ 4.7 Case Study
We present an example question from MuSiQue and compare the RAG process of NaiveRAG w/ QD and ChainRAG in Table 3. As NaiveRAG w/ QD does not rewrite the sub-question, its second sub-question lacks a clear entity, leading to retrieval errors and causing the “lost-in-retrieval” problems. This ultimately results in an incorrect answer. In the example, the context comes from the first chunk retrieved, which is also the source of the final answer. For ChainRAG, by using the answer from the first sub-question to complete the entity in the second sub-question, the retrieved information becomes more accurate, on which the LLM then provides the correct answer.
4.8 Further Analysis
Our method consists of several key steps, including entity recognition, question decomposition, and entity recovery. To assess their impact, we conduct additional experiments using Qwen2.5-72B.
4.8.1 Impact of Entity Recognition
Evaluating entity recognition performance on QA datasets is challenging since there is no established gold standard. To evaluate the robustness of ChainRAG in entity recognition, we integrate different entity recognition methods and compare their impact on the final QA performance. Specifically, we use Qwen2.5-72B for entity recognition and compare the QA F1 scores on MuSiQue against those obtained using spaCy. The results are presented in Table 4. We can see that the performance gains achieved with Qwen2.5-72B are modest. Our method is robust in entity extraction. Considering efficiency and effectiveness trade-offs, spaCy remains a well-balanced choice for entity extraction.
Method spaCy Qwen2.5-72B
Ours (AnsInt) 47.75 49.21
Ours (CxtInt) 49.37 50.61
Table 4: F1 scores on the MuSiQue dataset using different entity recognition methods with Qwen2.5-72B.
4.8.2 Impact of Question Decomposition and Entity Recovery
To evaluate question decomposition and entity recovery, we manually review the results of Qwen2.5-72B on the test examples (600 questions in total) of three datasets. On average, 89.3% of questions are successfully decomposed into reasonable sub-questions. Decomposition errors typically arise from questions containing lengthy modifiers or complex phrasing. In contrast, the accuracy of entity recovery is slightly lower, with an average score of 79.3%. This discrepancy can be attributed to the inherent difficulty of the entity recovery task, which requires LLMs not only to identify the demonstrative pronoun to be replaced but also to integrate the answer from the preceding sub-question to perform the replacement. Furthermore, to analyze the effect of question decomposition (QD) and entity recovery (ER) on the final QA performance, we divide the test examples into the following four groups:
  • Incorrect QD: The test examples that have incorrect question decomposition.
  • Incorrect ER: The test examples that have correct question decomposition but incorrect entity recovery results.
  • Others: The remaining test examples that have correct decomposition and entity recovery, or the examples do not require entity recovery (e.g., parallel multi-hop questions).
  • Total: The total test examples.
We calculate the F1 scores for each group using the sub-context integration method with Qwen2.5-72B. Despite incorrect question decomposition, the sub-context integration method demonstrates robustness, achieving an average F1 score of 55.43 by leveraging all retrieved information. This is often because one sub-question in an incorrect decomposition still resembles the original question, helping retrieval. However, incorrect entity recovery has a more significant negative impact, leading to “lost-in-retrieval” problems and lower performance compared to the “Others” group, which is unaffected by such problems. Overall, entity recovery errors are more detrimental to performance than question decomposition errors, underscoring the importance of addressing “lost-in-retrieval” problems. However, despite these problems, our method consistently outperforms the baselines, demonstrating its overall effectiveness and robustness.
MuSiQue 2Wiki HotpotQA
Incorrect QD 40.79 63.19 62.28
Incorrect ER 39.75 57.58 56.69
Others 54.30 67.12 65.83
Total 49.37 65.85 64.54
Table 5: F1 scores of different groups ChainRAG (CxtInt) with Qwen2.5-72B.
@@ Acknowledgments
This work was funded by National Natural Science Foundation of China (Nos. 62406136 and 62272219), Natural Science Foundation of Jiangsu Province (No. BK20241246), Postdoctoral Fellowship Program of CPSF (No. GZC20240685), China Postdoctoral Science Foundation (No. 2024M761396), and Jiangsu Funding Program for Excellent Postdoctoral Talent.
@@ Appendix C Detailed Results of Ablation Study
Tables 7 and 8 present the complete results of our ablation study. After removing certain processes from ChainRAG, we observe varying degrees of performance degradation in QA tasks. When the rewriting phase is removed, the performance on MuSiQue and HotpotQA drops significantly, while the performance on 2Wiki dataset sees a minor decrease. This is mainly because many of the questions in the 2Wiki dataset, after being decomposed, result in two parallel sub-questions with no dependency, which avoids the “lost-in-retrieval” problems. For MuSiQue, removing the SS edges has the greatest impact on performance, while for 2Wiki, it is the removal of the EC edges that has the most significant effect. In HotpotQA, the SA edges have the largest impact when removed. When we remove the entire graph and implement the indexing process according to NaiveRAG, we observe a noticeable decline in performance, with the CxtInt method being more significantly affected. This indicates the effectiveness of our method, which involves building an index using sentences and entities and performing retrieval on the sentence graph.
@@ Appendix D Prompt Example
Appendix E Additional Experiment on Small LLMs
To test the effectiveness of our method on small LLMs, we select two open-source small models, Qwen2.5-14B and Qwen2.5-7B, and compare their performance against the strongest two baselines from the main results. The results in Table 9 show that our method still outperforms the baselines, showcasing its effectiveness and reliability. These results also reflect the effectiveness of small LLMs in decomposition, rewriting, and summarization.
MuSiQue 2Wiki HotpotQA
14B LongRAG 36.59 54.95 58.90
Hipporag w/ IRCoT 37.09 58.50 54.51
Ours (AnsInt) 44.47 59.71 57.06
Ours (CxtInt) 43.16 63.14 59.43
7B LongRAG 29.01 49.02 54.71
Hipporag w/ IRCoT 29.30 51.17 53.79
Ours (AnsInt) 30.90 48.25 48.50
Ours (CxtInt) 33.15 52.18 60.03
Table 9: Performance of our ChainRAG using Qwen2.5-14B and 7B on MuSiQue, 2Wiki, and HotpotQA.
A. Examples of nodes
Node ID Corresponding sentence
Node1 Artificial intelligence has made tremendous progress in recent years.
Node2 Deep learning models can process vast amounts of data and identify patterns.
Node3 These models perform particularly well in image recognition tasks.
Node4 Natural language processing is an important branch of AI.
Node5 NLP enables machines to understand and generate human language.
Node6 This technology has been widely applied in virtual assistants.
Node7 Deep learning algorithms typically require large amounts of data.
Node8 Image recognition can help autonomous vehicles identify road signs.
Node9 Ethical considerations are an important factor in AI development.
Node10 Machine learning models learn from examples rather than explicit programming.
Node11 Neural networks are inspired by the structure of the human brain.
Node12 The development of AI may have significant impacts on the job market.
Node13 Reinforcement learning has enabled AI to master complex games.
Node14 AI systems are increasingly being used in healthcare for diagnosis.
Node15 Privacy concerns arise when AI systems collect and analyze personal data.
B. Examples of edges
Edge type Examples
SA edges (Node1, SA, Node2), (Node2, SA, Node3), (Node3, SA, Node4), (Node4, SA, Node5), …
SS edges (Node1, SS, Node13), (Node1, SS, Node12), (Node4, SS, Node1), (Node6, SS, Node1), (Node9, SS, Node15), …
EC edges (Node2, EC, Node3), (Node2, EC, Node10), (Node3, EC, Node10), (Node2, EC, Node7), …
Table 10: An example of our sentence graph.
Original question Sub-question 1 Rewriting Sub-question 2
Original question: What record label is the performer who released All Your Faded Things on? ChainRAG answer: Blue Note. Golden answer: Kill Rock Stars. First sub-question: Who is the performer of the song All Your Faded Things? Retrieved context: Personnel: Anna Oxygen - vocals, piano, composer, art design … LLM answer: Unable to answer the question. Correct answer: Anna Oxygen. Rewriting: Cannot rewrite, as the previous sub-question answer is “unable to answer”. Second sub-question: What record label is associated with this performer? Retrieved context: Jack Wilson featuring performances recorded and released on the Blue Note label… LLM answer:Blue Note. Correct answer:Kill Rock Stars.
Table 11: A failure case study.
Appendix F An Example of Sentence Graph
To help explain our sentence graph more clearly, we build an example using 15 sentences and present it in Table 10. SA edges connect directly adjacent sentences like (Node1, SA, Node2), (Node2, SA, Node3). For SS edges, such as (Node1, SS, Node13) and (Node1, SS, Node12), these connections indicate that Node13 and Node12 are the two sentences most semantically similar to Node1 within this graph. Other SS edges like (Node4, SS, Node1) and (Node6, SS, Node1) further illustrate these similarity-based connections. EC edges illustrate how sentences sharing key entities are linked. For instance, Node2, Node3, and Node10 are interconnected, forming edges like (Node2, EC, Node3), (Node2, EC, Node10), and (Node3, EC, Node10), because they all share the key entity “models”. Similarly, an EC edge exists between Node2 and Node7 since they both contain the key entity “data”.
Appendix G Failure Case Analysis
In this analysis, we provide a real error example from our test results to gain deeper insights into potential deficiencies that may constrain our method. From this example detailed in Table 11, we can observe several key issues. For the first sub-question, although the retrieved context contained the correct answer (“Anna Oxygen”), the LLM failed to provide it and instead responded with “Unable to answer”. We attribute this to potential limitations in the LLM’s ability to process long contexts effectively. This failure meant that the required entity for the second sub-question was unavailable, leading to a “lost-in-retrieval” problem. Consequently, the retrieval process for the second sub-question failed to obtain relevant contextual information about “Anna Oxygen”. This chain of errors ultimately led to the incorrect answer (“Blue Note”).